Seismicity during the past 24 hours has been of a variable nature with long period earthquakes and rockfalls still the most dominant feature. Numerous long period earthquakes, which are too small to locate, occurred at a rate of approximately one per minute throughout the review period. The low to moderate level broadband tremor, observed at 16:30 yesterday, decreased in amplitude by approximately 20:00 (29 February) and resumed by 11:00 (01 March). Numerous volcano-tectonic earthquakes were detected on the seismograph stations nearest the volcano, but were too small to be located. Seismic signals, associated with rock fall events, were detected throughout the review period, but generally the overall level of activity was lower than the previous 24 hour period.
EDM measurements were done at the eastern, western and southern triangles of the volcano today under good viewing conditions and once again there were no significant changes between the measured lines.
Visual observations were made from the observation point in Farrell's in the late afternoon. Although no major rockfalls were seen by the observers at Farrell's there was clear evidence that debris was continuing to fill the western and northern moat areas of the crater. The volcano was still venting generously from the central and eastern sections of the dome. A helicopter flight yesterday evening (29 February) revealed several areas of intense glowing, and some small avalanches of glowing rock were noted from the central and northwestern parts of the dome.
Mr Angus Miller, from the BGS, returned to the observatory today to replace Mr Lutchman Pollard of SRU, whose duties were in the area of seismic data processing and interpretation.